OpenStudy (anonymous):

2x+3=5 help needed in x @chosenmatt

OpenStudy (anonymous):

x=1

OpenStudy (chosenmatt):

x=1

OpenStudy (anonymous):

because 2 times 1+3 =5

OpenStudy (anonymous):

2x+3=5 Subtract 3 from both sides 2x=2 divide 2 by both sides x=1

OpenStudy (chosenmatt):

yup :)

OpenStudy (anonymous):

2x+3=5 Subtract 3 from both sides 2x=2 divide 2 by both sides x=1

OpenStudy (anonymous):

did this help @surd ?